the pressure of oxygen gas in an elastic balloon at 25C is 1 atm. What is the pressure in atm after the temperature is icreased to 37C


1
Expert's answer
2021-02-12T05:03:59-0500

"\frac{P_1}{T_1}" = "\frac{P_2}{T_2}"


"\frac{1atm}{25⁰C}" = "\frac{P_2}{37⁰C}"


P2 = 1.48 atm